Kunkel Company makes two products and uses a conventional costing system in which a single plantwide predetermined overhead rate is computed based on direct labor-hours. Data for the two products for the upcoming year follow:

Mercon Wurcon
Direct materials cost per unit $ 10.00 $ 8.00
Direct labor cost per unit $ 3.00 $ 3.75
Direct labor-hours per unit 0.20 0.25
Number of units produced 10,000 40,000

These products are customized to some degree for specific customers.

Required:

1. The company's manufacturing overhead costs for the year are expected to be $336,000. Using the company's conventional costing system, compute the unit product costs for the two products.

2. Management is considering an activity-based costing system in which half of the overhead would continue to be allocated on the basis of direct labor-hours and half would be allocated on the basis of engineering design time. This time is expected to be distributed as follows during the upcoming year:

Mercon Wurcon Total
Engineering design time (in hours) 4,000 4,000 8,000

Compute the unit product costs for the two products using the proposed ABC system.

(For all requirements, do not round intermediate calculations. Round your final answers to 2 decimal places.)
Answer #1
1. Computation of unit product costs for the two products using company's conventional costing system:-
A. Computation of Single Plantwide Manufacturing overhead rate :-
Estimated Total manufacturing Overheads $             336,000
Estimated total Direct labor hours
Product Direct labor-hours per unit Number of units produced Estimated Total Direct labor-hours
a b c d = b*c
Mercon 0.20                     10,000                      2,000 Direct labor-hours
Wurcon 0.25                     40,000                   10,000 Direct labor-hours
Estimated total Direct labor hours                   12,000 Direct labor-hours
Single Plantwide Manufacturing overhead rate = Estimated Total manufacturing Overheads
Estimated total Direct labor hours
= $336,000
12000
= $28.00 per Direct labor-hour
B. Computation of Overhead Cost Per unit:-
Product Activity Driver Plantwide OH Rate Overhead Cost
a d = b*c c d = b×c
Mercon                      2,000 Direct labor-hours $28.00 per Direct labor-hour $                        56,000
Wurcon                    10,000 Direct labor-hours $28.00 per Direct labor-hour $                      280,000
C. Computation of Unit product costs :-
Product
Mercon Wurcon
No. of units produced and sold 10000 40000
Per Unit Total Per Unit Total
Direct Material Cost $                   10.00 $             100,000 $                                8.00 $                      320,000
Direct labor Cost $                     3.00 $               30,000 $                                3.75 $                      150,000
Overhead Cost $                     5.60 $               56,000 $                                7.00 $                      280,000
Unit product costs $                   18.60 $             186,000 $                              18.75 $                      750,000
2. Computation of unit product costs for the two products using activity based costing system:-
A. Computation of Cost per driver :-
Activity cost pool Activity measures (Cost Drivers) Estimated Overhead Cost Expected Activity Cost per driver
Mercon Wurcon Total
a b c d e f g = c/f
Cost Pool 1 (half of $336000) Direct labor-hours $              168,000                        2,000                   10,000                                12,000 $                          14.00
Cost Pool 2 (half of $336000) Engineering design hours $              168,000                        4,000                      4,000                                   8,000 $                          21.00
Total Manufacturing Overhead Cost $             336,000
B. Computation of Overhead Cost Per unit:-
Overhead Assigned Activity Driver Cost per driver Total Overhead Cost
a b c d = b×c
Mercon
Cost Pool 1 (half of $336000)                      2,000 Direct labor-hours $                              14.00 $                        28,000
Cost Pool 2 (half of $336000)                      4,000 Engineering design hours $                              21.00 $                        84,000
Total $                      112,000
Units Produced                            10,000
Overhead Cost per Unit $                          11.20
Wurcon
Cost Pool 1 (half of $336000)                    10,000 Direct labor-hours $                              14.00 $                      140,000
Cost Pool 2 (half of $336000)                      4,000 Engineering design hours $                              21.00 $                        84,000
Total $                      224,000
Units Produced                            40,000
Overhead Cost per Unit $                             5.60
C. Computation of Unit product costs:-
Product
Mercon Wurcon
No. of units produced 10000 40000
Per Unit Total Per Unit Total
Direct Material Cost $                   10.00 $             100,000 $                                8.00 $                      320,000
Direct labor Cost $                     3.00 $               30,000 $                                3.75 $                      150,000
Overhead Cost $                   11.20 $             112,000 $                                5.60 $                      224,000
Unit product costs $                   24.20 $             242,000 $                              17.35 $                      694,000

Feel free to ask any clarification, if required. Please provide feedback by thumbs up, if satisfied. It will be highly appreciated. Thank you.